OpenStudy (anonymous):

A Line has a slope of -2/3 and passes throug the points (-3, 8). what is the equation of the line? @AriPotta

OpenStudy (aripotta):

so the way to go with this is to plug your info into point-slope form\[y - y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

ok so what would be the y, the m and the x?

OpenStudy (aripotta):

you have a point (-3,8). a point is in the form (x,y) so -3 will be your x1 and 8 will your y1. the y and x stay the same and the m is your slope, -2/3

OpenStudy (aripotta):

y - 8 = -2/3(x + 3)

OpenStudy (aripotta):

does that make sense?

OpenStudy (anonymous):

yea that makes sense but the answers choices are different on mine

OpenStudy (aripotta):

can you post the choices?

OpenStudy (aripotta):

they probably want you to write it in a different form

OpenStudy (anonymous):

y=-2/3x+6 y=-2/3x+8 y=6x-2/3 y=8x-2/3

OpenStudy (aripotta):

yea, so those choices are in slope-intercept form. we can easily convert our equation into slope-intercept

OpenStudy (anonymous):

ok, how do we do that?

OpenStudy (aripotta):

y - 8 = -2/3(x + 3) ok. so first distribute the -2/3

OpenStudy (anonymous):

how do i distribute them? by dividing?

OpenStudy (aripotta):

multiplying. so -2/3 times x = -2/3x. -2/3 times 3 = -2 y - 8 = -2/3x - 2

OpenStudy (aripotta):

then, to get y by itself, you need to add 8 to both sides

OpenStudy (anonymous):

y=-2/3x+6?

OpenStudy (aripotta):

yep

OpenStudy (anonymous):

ok thanks!

OpenStudy (aripotta):

no problemo :)